TOWARD A COMPLETE EQUILIBRIUM STRUCTURE OF BUTADIENE; HIGH-RESOLUTION IN-FRARED SPECTROSCOPY OF BUTADIENE-1-^{13}C_{1}

\underline{NORMAN C. CRAIG}, KEITH A. HANSON, and MICHAEL C. MOORE, Department of Chemistry, Oberlin College, Oberlin, OH 44074; ROBERT L. SAMS, Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352.

Considerable progress has been made toward obtaining ground state rotational constants for butadiene (BDE) and its isotopomers for use in fitting an equilibrium structure. With the exception of a microwave investigation of the weakly polar BDE-1,1-d_{2},^{a} studies of all of the other, nonpolar species have been done with high-resolution (0.002 cm^{−1}) infrared spectroscopy. Rotational constants are available for BDE and BDE-2,3-d_{2} from one study^{b} and for the three species of BDE-1,4-d_{2} from another study.^{c} The present report is on BDE-1-^{13}C_{1}. The rotational structure in the C-type bands at 524 cm^{−1}, 900 cm^{−1}, and 908 cm^{−1} in the infrared spectrum has been analyzed. Rotational constants fit to 2191 ground state combination differences derived from all three bands are (in cm^{−1}) A = 1.3887919 (6), B = 0.1436683 (3), and C = 0.1302251 (3). In the parent molecule of C_{2h} symmetry, a Raman-active b_{g} mode occurs at 908 cm^{−1} and an infrared-active mode occurs at essentially the same frequency. In BDE-1-^{13}C_{1} of reduced, C_{s} symmetry, both modes have significant infrared intensity and occur at 908 and 900 cm^{−1}. The higher frequency mode is CH_{2} flapping; the lower frequency one is ^{13}CH_{2} flapping. ^{d}

aW. Caminati, G. Grassi, and A. Bauder Chem. Phys. Letters 148, 13, (1988).

bN. C. Craig, J. L. Davis, K. A. Hanson, M. C. Moore, and M. Lock J. Mol. Struct. 00, 000, (2004).

cN. C. Craig, K. A. Hanson, R. W. Pierce, S. D. Saylor, and R. L. Sams J. Mol. Spectrosc., submitted.

dP. Huber-Wälchli and Hs. H. Günthard Spectrochim. Acta, 37A, 285, (1981); Yu. N. Panchenko, private communication.